Shift Summary for Owl, Feb. 2, 2000


Subject: Shift Summary for Owl, Feb. 2, 2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:49:46 CST


              Shift Summary for owl shift, 02/Feb/2000

On Shift: S. Ledovskoy, P. Gouffon + BobH & R.Kessler

Hours of beam delivered at all intensities: 6.0/7.0
Hours of KAMI data: 5.5
Accounting of lost time: Booster problems, run change

========================================================================
Runs ended this shift:
  Run 15737 - KAMI_KE3
  Run 15738 - KAMI_MUON
  Run 15739 - KAMI_MUON

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
  Continuation of KAMI data collection.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
              00:00 Take over from evening shift. BobT and scottish
                        dancer D.Jensen present, to leave soon.
15737 01:00 Run started before but no beam
              01:09 beam is back
              02:40 run ended

              02:40 switching to muon
15738 1 02:45 started run 15738, KAMI_MUON
              02:50 run stopped: B06 not on process list

15739 1 03:00 started run 15739, KAMI_MUON
              04:00 gas check list done. BIC softfeel pen doesn't
                        work in cold weather (5C in gas shed)
              04:10 RK leaves
              04:20 BobH leaves
         143 06:15 End of run 15739. 808K events
                        MCR called about access to NM2 to move to
                        15mrad targetting angle, which was scheduled for 0600.
              06:55 nobody showed up yet for that job!
              07:00 cleaning crew changed black carpet
========================================================================



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:49:49 CST